# DIRESA, a distance-preserving nonlinear dimension reduction technique based on regularized autoencoders

### v1

@@ 4.1 Lorenz ’63
In our benchmark, we use nine different DR techniques to compress the 3-dimensional Lorenz butterfly into a 2-dimensional latent space: Principal Component Analysis (PCA), kernel PCA (KPCA), Uniform Manifold Approximation and Projection (UMAP), and six artificial neural networks (ANN): AE, AE with batch normalization on the latent space (BNAE), AE with covariance regularization on the latent space (CRAE), variational AE (VAE) and two distance regularized siamese twin AE (DIRESA).
The hyperparameters for the nine different DR techniques are listed in Table LABEL:table:ModelConfLorenz. Hyperparameter tuning is done using simulated annealing [kirkpatrick1983optimization] for KernelPCA and UMAP. More information on the hyperparameter tuning can be found in the Supporting Information. All ANN methods have an encoder submodel with a 3-dimensional input layer, two hidden layers with 40 and 20 fully connected nodes, and an output layer with two fully connected nodes, see Table LABEL:table:ModelDiresaLorenz.
@@ 4.1.1 Evaluation of latent and decoded samples
The different DR methods are evaluated quantitatively based on the loss function output of the test dataset (see Table 1). The properties of the latent space are assessed qualitatively by visualizing the decoding uniform random points in latent space (see Figure 3) and by visualizing the encoded test dataset in latent space (see Figure 4). In terms of reconstruction loss (or unexplained variance), PCA performs the worst (because of the linearity), and the latent space is decoded into a flat plane (see Figure 3(a)). A standard AE has the lowest reconstruction loss (the lowest unexplained variance). Still, the decoded random sample shown in Figure 3(d) is irregular and fails to generalize well outside the wings. BNAE, CRAE, and VAE are also doing well from a reconstruction loss point of view. Both DIRESA methods are doing better for the reconstruction loss than KPCA and UMAP, although a trade-off had to be made between the three different loss functions. The decoded latent space curves for both DIRESA variants follow the wings, and outside the wings, the curve is well generalized (see Figure 3(h) and 3(i)), which is also the case for BNAE (see Figure 3(e)) and VAE (see Figure 3(g)). Using an MSE or a Corr distance loss function makes little difference from a reconstruction point of view. Unlike the other methods, the UMAP decoded sample is more scattered (see Figure 3(c)) and does not lie in a 2-dimensional manifold.
Figure 4 shows the mapping of the Lorenz test dataset into latent space. The encoded latent components are uncorrelated for all DR methods except for the standard AE (see Figure 4(d)). As we can see in Figure 4(a) and 4(b), PCA and KPCA order the latent components by variance, which is not the case for the different ANNs, see Figure 4(g) for the VAE and Figure 4(h) for DIRESA, where the wing is rotated. The wing structure is well preserved with PCA, KPCA, VAE, and DIRESA methods, which is less the case for UMAP (see Figure 4(c)) and the AE. With a standard AE, no regularization is done on latent space, so we get different irregular pictures for other training runs. For BNAE, both latent components have a standard deviation of 1, so the wing structure is resized in one direction (see Figure 4(e)). CRAE has no distance regularization, so one wing is more prominent here than the other (see Figure 4(f)).
@@ 4.1.2 Distance ordering preservation
DIRESAMSE 0.997 0.997 0.107 0.849 0.896 0.848
DIRESACorr 0.997 0.997 0.112 0.845 0.894 0.843
Table 2: Mean distance ordering preservation between original and latent space for the Lorenz ’63 test dataset. The Corr and LogCorr columns show the Pearson correlation for Euclidean and logarithmic distances. The Can50 column gives the Canberra stability indication (smaller is better), the Pear50, Spear50 and Ken50 show the Pearson, Spearman, and Kendall correlation with a location parameter of 50. The 3 best values are shaded in red; the best is dark red.
Table 2 shows how distance ordering is preserved between original and latent space for the different DR methods. We calculated the distances from all points of the test dataset to all other points and did the same for the corresponding points in latent space. The Corr column shows the average Pearson correlation between the Euclidean distances in original and latent space. The LogCorr shows the same but for the logarithm of the distances. When searching for analogs, we are only interested in the closest distances. In the right four columns of Table 2, a location parameter 50 has been set, meaning that only the 50 closest distances (in latent space) are considered. The Can50 shows the mean Canberra stability indicator between the Euclidean distances. The Canberra stability indicator measures the difference between 2 ranked lists (smaller is better). A Canberra stability indicator of about 0 means the lists are ranked very similarly, while two random lists will result in a figure of around 1.42. Rank differences at the top of the list are more penalized than differences in the bottom part. The Pear50, Spear50 and Ken50 columns are showing Pearson’s r, Spearman’s $\rho$ and Kendall’s $\tau$ correlation with a location parameter of 50 (in the Supporting Information Table LABEL:table:LorenzDistPreservationMean100 shows the same indices for a location parameter of 100). Table 3 shows the median of the same indicators. The three best scores are shown in red, and the best score is in dark red. The Supporting Information Table LABEL:table:LorenzDistPreservationStdErr50 shows the standard error of the mean for the different indicators. Table LABEL:table:LorenzDistPreservationPvalue shows the 2 sample t-test p-value for the mean distance between DIRESA and PCA. Figure 5 shows the scatter plots of the 50 closest distances (in latent space) for the different methods for 200 random samples of the test dataset. The points in the two wings’ cross-section are where the distance ordering preservation is worst when mapped on a 2-dimensional curve.
@@
DIRESAMSE 0.999 0.999 0.040 0.995 0.990 0.935
DIRESACorr 0.999 0.999 0.045 0.994 0.989 0.928
Table 3: Median distance ordering preservation between original and latent space for the Lorenz ’63 test dataset. Columns and shading as in Table 2.
Figure 5: Scatter plots of 50 closest distances for 200 random samples of the test dataset for PCA (5(a)), KPCA (5(b)), UMAP (5(c)), AE (5(d)), BNAE (5(e)), CRAE (5(f)), VAE (5(g)), DIRESAMSE (5(h)) and DIRESACorr (5(i))
@@ 4.2 MAOOAM
Figure 7: Pair plot for MAOOAM latent components for PCA (7(a))
and DIRESAMSE (7(b))
@@
Figure 8: Time plot for MAOOAM latent components for PCA (8(a))
and DIRESAMSE (8(b))
For MAOOAM, we will compare DIRESA with PCA. The complete DIRESA configuration can be found in the supporting information. The encoder of the DIRESA DR methods has one convolutional layer, followed by a MaxPooling and three fully connected layers, see Table LABEL:table:ModelDiresaMA. The values for all other hyperparameters can be found in LABEL:table:ModelConfMA.
@@ 4.2.1 Evaluation of latent and decoded samples
Figure 9: MAOOAM: PCA decoded latent components
Figure 10: MAOOAM: DIRESA decoded latent components
Figure 7(b) shows the pair plots for the latent components of the test dataset for PCA and DIRESA with MSE distance loss (after sorting by explained variance). The DIRESA pair plots look very similar to those of PCA. Figure 8(b) shows the time plots for the latent components of the test dataset for PCA and DIRESA with MSE distance loss. Again, the DIRESA time plot looks very similar to the PCA’s. The first four latent components with the largest explained variance have a low-frequency pattern, while the last six latent components have a high frequency. Figure 9 shows the empirical orthogonal functions (or PCA eigenvectors) for the different variables together with the explained variance (overall and per variable). To get the decoded latent components for DIRESA, all latent components of the encoded test dataset are set to their mean except for one component, which is set firstly to the mean plus one standard deviation and secondly to the mean minus one standard deviation. Both are then decoded, and the difference is taken. The result is shown in Figure 10. Decoded components for a variable with an explained variance of less than $1.5$ are faded. Again, we notice that the decoded latent DIRESA components (with MSE distance loss) are very similar to the PCA empirical orthogonal functions, except that the explained variance is higher. The unexplained variance for PCA is 3.7%, and for DIRESA 2.7%, the difference is the biggest for the atmospheric temperature, where the unexplained variance for PCA is 9.3% and for DIRESA 4.1%. We observe a strong coupling between atmospheric and ocean variables for the low-frequency components. More than 95% of the ocean stream function variance is explained by the first two components and for the ocean temperature by the first four. There is no coupling between atmospheric and ocean variables for the high-frequency components, except for component 9, which couples the atmospheric stream function with the ocean temperature.
@@
PCA 0.999 0.998 0.570 0.432 0.398 0.292
DIRESAMSE 0.999 0.999 0.565 0.429 0.401 0.292
DIRESAMSLE 0.999 0.998 0.518 0.490 0.453 0.338
DIRESACorr 0.999 0.999 0.553 0.453 0.417 0.304
DIRESALogCorr 0.999 0.999 0.534 0.472 0.443 0.328
Table 4: Mean distance ordering preservation between original and latent space for 200 random samples of the MAOOAM test dataset. Columns and shading as in Table 2.
@@ 4.2.2 Distance ordering preservation
Table 4 shows the distance (ordering) preservation for PCA and DIRESA with four different distance loss functions: MSE, MSLE, correlation, and correlation of the logarithm of the distances. As it is computationally expensive to calculate distances from all points of the test dataset to all other points, we took 200 random samples from the test dataset and calculated the distance to all other points. 200 random samples are enough to distinguish the best DIRESA methods (in this case, DIRESAMSLE and DIRESALogCorr) from PCA with a good p-value (see Supporting Information Table LABEL:table:MADistPreservationPvalue). The figures are shown for the same KPIs as in paragraph 4.1.2. Looking at all the distances, near and far, PCA and DIRESA methods are performing equally for the Corr KPI, and three out of the four DIRESA methods are giving slightly better figures for the LogCorr KPI (which provides more emphasis on the short distances). Looking at the 50 closest distances, all DIRESA methods outperform PCA for Can50 and Spear50, and three out of four DIRESA methods outperform PCA for the Pear50 and Ken50 KPIs. Methods with a logarithmic distance loss function yield the best results.
@@ 4.3 Conclusion
In addition to the present study, we provide an open-source Python package, based on Tensorflow [tensorflow2015-whitepaper], to build DIRESA models with convolutional and/or dense layers with one line of code [depaepe2024diresa]. On top of that, the package allows the use of custom encoder and decoder submodels to build a DIRESA model. The DIRESA package acts as a meta-model, which can use submodels with various kinds of layers, such as attention layers, and more complicated designs, such as graph neural networks or hybrid quantum neural networks. Thanks to its extensible design, the DIRESA framework can handle more complex data types, such as three-dimensional, graph, or unstructured data.
Its flexibility and robust performance make DIRESA an indispensable new tool in weather and climate science. It has all the assets to become the method of choice to distil meaningful low-dimensional representations from the ever-increasing volumes of high-resolution climate data, for applications ranging from analog retrieval to attribution studies.
@@ 5 Data availability
The data used in this study is available on the open data portal Zenodo [depaepe2024dataset].
@@ 6 Code availability
The DIRESA package source code is available on the open data portal Zenodo [depaepe2024diresa]. Other programs and scripts are available from the corresponding author upon reasonable request. Documentation can be consulted at https://diresa-learn.readthedocs.io.

### v2

@@ 4.1 Lorenz ’63
In our benchmark, we use nine different DR techniques to compress the 3-dimensional Lorenz butterfly into a 2-dimensional latent space: Principal Component Analysis (PCA), kernel PCA (KPCA), Uniform Manifold Approximation and Projection (UMAP), and six ANNs: AE, AE with batch normalization on the latent space (BNAE), AE with covariance regularization on the latent space (CRAE), variational AE (VAE) and two distance regularized siamese twin AE (DIRESA).
The hyperparameters for the nine different DR techniques are listed in Table 6. Hyperparameter tuning is done using simulated annealing (Kirkpatrick et al., 1983) for KernelPCA and UMAP. More information on the hyperparameter tuning can be found in Appendix B. All ANN methods have an encoder submodel with a 3-dimensional input layer, two hidden layers with 40 and 20 fully connected nodes, and an output layer with two fully connected nodes, see Table 7. This results in an AE and DIRESA ANN with 2045 trainable parameters.
@@ 4.1.1 Evaluation of latent and decoded samples
The different DR methods are evaluated quantitatively based on the loss function output of the test dataset (see Table 1). The properties of the latent space are assessed qualitatively by visualizing the decoding uniform random points in latent space (see Figure 3) and by visualizing the encoded test dataset in latent space (see Figure 4). In terms of reconstruction loss, PCA performs the worst (because of the linearity), and the latent space is decoded into a flat plane (see Figure 3a). A standard AE has the lowest reconstruction loss (and so the highest $R^{2}$ score). Still, the decoded random sample shown in Figure 3d is irregular and fails to generalize well outside the wings. BNAE, CRAE, and VAE are also doing well from a reconstruction loss point of view. Both DIRESA methods are doing better for the reconstruction loss than KPCA and UMAP, although a trade-off had to be made between the three different loss functions. The decoded latent space curves for both DIRESA variants follow the wings, and outside the wings, the curve is well generalized (see Figure 3h and 3i), which is also the case for BNAE (see Figure 3e) and VAE (see Figure 3g). Using an MSE or a Corr distance loss function makes little difference from a reconstruction point of view. Unlike the other methods, the UMAP decoded sample is more scattered (see Figure 3c) and does not lie in a 2-dimensional manifold.
Figure 4 shows the mapping of the Lorenz test dataset into latent space. The encoded latent components are uncorrelated for all DR methods except for the standard AE (see Figure 4d). As we can see in Figure 4a and 4b, PCA and KPCA order the latent components by variance, which is not the case for the different ANNs, see Figure 4g for the VAE and Figure 4h for DIRESA, where the wing is rotated. The wing structure is well preserved with PCA, KPCA, VAE, and DIRESA methods, which is less the case for UMAP (see Figure 4c) and the AE. With a standard AE, no regularization is done on latent space, so we get different irregular pictures for other training runs. For BNAE, both latent components have a standard deviation of 1, so the wing structure is resized in one direction (see Figure 4e). CRAE has no distance regularization, so one wing is more prominent here than the other (see Figure 4f).
@@ 4.1.2 Distance ordering preservation
DIRESAMSE 0.997* 0.997* 0.107* 0.849 0.896* 0.848*
DIRESACorr 0.997* 0.997* 0.112 0.845 0.894 0.843
Table 2: Mean distance ordering preservation between original and latent space for the Lorenz ’63 test dataset. The Corr and LogCorr columns show the Pearson correlation for Euclidean and logarithmic distances. The Can50 column gives the Canberra stability indication (smaller is better), the Pear50, Spear50 and Ken50 show the Pearson, Spearman, and Kendall correlation with a location parameter of 50. The 3 best values are highlighted in bold; the best one is marked with an asterisk.
Table 2 shows how distance ordering is preserved between original and latent space for the different DR methods. We calculated the distances from all points of the test dataset to all other points and did the same for the corresponding points in latent space. The Corr column shows the average Pearson correlation between the Euclidean distances in original and latent space. The LogCorr shows the same but for the logarithm of the distances. When searching for analogs, we are only interested in the closest distances. In the right four columns of Table 2, a location parameter 50 has been set, meaning that only the 50 closest distances (in latent space) are considered. The Can50 shows the mean Canberra stability indicator between the Euclidean distances. The Canberra stability indicator measures the difference between 2 ranked lists (smaller is better). A Canberra stability indicator of about 0 means the lists are ranked very similarly, while two random lists will result in a value of around 1.42. Rank differences at the top of the list are more penalized than differences in the bottom part. The Pear50, Spear50 and Ken50 columns are showing Pearson’s r, Spearman’s $\rho$ and Kendall’s $\tau$ correlation with a location parameter of 50 ( Appendix Table 12 shows the same indices for a location parameter of 100). Table 3 shows the median of the same indicators. The three best scores are shown in red, and the best score is in dark red. The Appendix Table 10 shows the standard error of the mean for the different indicators. Table 11 shows the 2 sample t-test p-value for the mean distance between DIRESA and PCA. Figure 5 shows the scatter plots of the 50 closest distances (in latent space) for the different methods for 200 random samples of the test dataset. The points in the two wings’ cross-section are where the distance ordering preservation is worst when mapped on a 2-dimensional curve.
@@
DIRESAMSE 0.999* 0.999* 0.040* 0.995* 0.990* 0.935*
DIRESACorr 0.999* 0.999* 0.045 0.994 0.989 0.928
Table 3: Median distance ordering preservation between original and latent space for the Lorenz ’63 test dataset. Columns and highlights as in Table 2.
Figure 5: Scatter plots of 50 closest distances for 200 random samples of the Lorenz test dataset for for the different DR techniques
@@ 4.2 MAOOAM
Figure 7: Pair plot for MAOOAM latent components for PCA (a)
and DIRESAMSE (b)
For MAOOAM, we will compare DIRESA with PCA. The complete DIRESA configuration can be found in Appendix B. The encoder of the DIRESA DR methods has one convolutional layer, followed by a MaxPooling and three fully connected layers, see Table 8. This results in a DIRESA ANN with 791,314 trainable parameters. The values for all other hyperparameters can be found in 9.
Figure 8: Time plot for MAOOAM latent components for PCA (a)
and DIRESAMSE (b)
@@ 4.2.1 Evaluation of latent and decoded samples
Figure 9: MAOOAM: PCA empirical orthogonal functions for the atmospheric stream function, atmospheric temperature, ocean stream function and ocean temperature. Figures show the total explained variance (on the left) and the explained variance per variable. Decoded components for a variable with an explained variance of less than 1.5 are faded.)
Figure 10: MAOOAM: DIRESA decoded latent components for the atmospheric stream function, atmospheric temperature, ocean stream function and ocean temperature. Figures show the total $R^{2}$ score (on the left) and the $R^{2}$ score per variable. Decoded components for a variable with an $R^{2}$ score of less than 1.5 are faded.)
Figure 7 shows the pair plots for the latent components of the test dataset for PCA and DIRESA with MSE distance loss (after sorting). The DIRESA pair plots look very similar to those of PCA. Figure 8 shows the time plots for the latent components of the test dataset for PCA and DIRESA with MSE distance loss. Again, the DIRESA time plot looks very similar to the PCA’s. The first four latent components with the largest $R^{2}$ score have a low-frequency pattern, while the last six latent components have a high frequency. Figure 9 shows the empirical orthogonal functions (or PCA eigenvectors) for the different variables together with the percentage of explained variance (overall and per variable). The decoded latent components for DIRESA are shown in Figure 10 (the procedure for calculating these is explained in Appendix D). Decoded components for a variable with an $R^{2}$ score of less than $1.5$ are faded. Again, we notice that the decoded latent DIRESA components (with MSE distance loss) are very similar to the PCA empirical orthogonal functions, except that the $R^{2}$ score is higher. The FVU for PCA is 3.7%, and for DIRESA 2.8%, the difference is the biggest for the atmospheric temperature, where the FVU for PCA is 9.3% and for DIRESA 6.0%. PCA shows a strong coupling between atmospheric and ocean variables for the first, third and fourth component (with an $R^{2}$ score of more than 5% for at least one of the atmospheric and ocean variables). A small coupling is shown for the second components, with an $R^{2}$ of 2% for the atmospheric temperature, 48% for the ocean stream function and 8% for the ocean temperature. DIRESA shows only a strong coupling for the third component and a low coupling for the fourth component. The first two components have a cumulative $R^{2}$ score (see Appendix D) of more than 95% for the ocean stream function (for both PCA and DIRESA). For the ocean temperature, the first four components are needed for a similar cumulative $R^{2}$ score. DIRESA shows no coupling between atmospheric and ocean variables for the high-frequency components, similarly to PCA.
@@
PCA 0.999* 0.998 0.570 0.432 0.398 0.292
DIRESAMSE 0.999* 0.999* 0.565 0.429 0.401 0.292
DIRESAMSLE 0.999* 0.998 0.518* 0.490* 0.453* 0.338*
DIRESACorr 0.999* 0.999* 0.553 0.453 0.417 0.304
DIRESALogCorr 0.999* 0.999* 0.534 0.472 0.443 0.328
Table 4: Mean distance ordering preservation between original and latent space for 200 random samples of the MAOOAM test dataset. Columns and highlights as in Table 2.
@@ 4.2.2 Distance ordering preservation
Table 4 shows the distance (ordering) preservation for PCA and DIRESA with four different distance loss functions: MSE, MSLE, correlation, and correlation of the logarithm of the distances. As it is computationally expensive to calculate distances from all points of the test dataset to all other points, we took 200 random samples from the test dataset and calculated the distance to all other points. 200 random samples are enough to distinguish the best DIRESA methods (in this case, DIRESAMSLE and DIRESALogCorr) from PCA with a good p-value (see Appendix Table 16). The figures are shown for the same KPIs as in paragraph 4.1.2. Looking at all the distances, near and far, PCA and DIRESA methods are performing equally for the Corr KPI, and three out of the four DIRESA methods are giving slightly better figures for the LogCorr KPI (which provides more emphasis on the short distances). Looking at the 50 closest distances, all DIRESA methods outperform PCA for Can50 and Spear50, and three out of four DIRESA methods outperform PCA for the Pear50 and Ken50 KPIs. Methods with a logarithmic distance loss function yield the best results.
@@ 4.3 Conclusion
In addition to the present study, we provide an open-source Python package, based on Tensorflow (Abadi et al., 2015), to build DIRESA models with convolutional and/or dense layers with one line of code (De Paepe, 2024b). On top of that, the package allows the use of custom encoder and decoder submodels to build a DIRESA model. The DIRESA package acts as a meta-model, which can use submodels with various kinds of layers, such as attention layers, and more complicated designs, such as graph neural networks or hybrid quantum neural networks. Thanks to its extensible design, the DIRESA framework can handle more complex data types, such as three-dimensional, graph, or unstructured data.
In the future, we will use the DIRESA DR method on real-world datasets such as weather radar data or reanalysis data. Thanks to the flexibility of the neural encoder, we expect the highest benefit compared to PCA and other traditional DR methods on datasets originating from systems with an attractor spanning a strongly nonlinear, lower-dimensional subspace. Large-scale atmospheric dynamics, quantified by daily sea level pressure fields, are an example of such a system (Faranda et al., 2017). For a purely stochastic system in which the attractor does not span a lower-dimensional subspace, we do not expect DIRESA to perform better than other DR methods. A prerequisite for using DIRESA is that the training dataset is large enough, which is less necessary for PCA.
We believe that DIRESA’s flexibility and robust performance give it all the assets to become a useful tool to distil meaningful low-dimensional representations from the ever-increasing volumes of high-resolution climate data.
@@ 5 Data availability
The data used in this study is available on the open data portal Zenodo (De Paepe, 2024a).
@@ 6 Code availability
The DIRESA package source code is available on the open data portal Zenodo (De Paepe, 2024b). Other programs and scripts are available from the corresponding author upon reasonable request. Documentation can be consulted at https://diresa-learn.readthedocs.io.